# Precise Constraints on the Energy Budget of WASP-121 b from its JWST NIRISS/SOSS Phase Curve

### v1

@@ 2.3.3 Gaussian Process
We implement a Gaussian Process (GP) model to marginalize our inferences over leftover non-white noise. We use the open-source Python package celerite (Foreman-Mackey et al., 2017b), and consider the SHOTerm Kernel which represents a stochastically driven, damped harmonic oscillator. This kernel models the power spectral density, , using three input parameters: the signal amplitude , characteristic frequency , and quality factor :
@@
The quality factor is set to , a commonly used value for modeling stellar granulation. We use the scaling relations of Gilliland et al. (2011) (equations 5 & 6 of that work) and find an expected granulation signal timescale of 6 minutes and an amplitude of 96 ppm, a level we should be sensitive to. Following Radica et al. (2024) and Coulombe et al. (2025), we adopt the parameterization of Pereira et al. (2019) and reframe the GP parameters in terms of amplitude, , and timescale, . These are then converted back to and before being passed to celerite.
2.3.4 Light Curve Fitting
@@
We fit our light curves using the emcee Python package, which implements a Monte Carlo ensemble sampler with affine-invariant Markov chains (Foreman-Mackey et al., 2013). We construct the white light curves by summing the flux across all 2048 pixels [0.85–2.85 m] for Order 1 and from [0.6–0.85 m] for Order 2. We then normalize both white light curves to the first eclipse and are then analyzed separately, with the orbital eccentricity fixed to zero (Delrez et al., 2016; Bourrier et al., 2020a).
@@
For each order, we fit a total of 21–23 parameters. The orbital parameters include the orbital period (), the semi-major axis relative to the stellar radius (), the mid-transit time (), and the orbital inclination (). The astrophysical parameters consist of the planet-to-star radius ratio (), the sub-stellar planet radius to stellar radius ratio (), the quadratic limb-darkening coefficients () (Kipping, 2013), and the eclipse depth (). To account for phase variations, we fit four sinusoidal parameters (). To avoid unphysical solutions, we impose a prior on the planetary flux (equation 4). Specifically, for any modeled point that yields negative flux, we assign a half-Gaussian prior centered at zero, with the width equal to the measurement uncertainty of that point. This construction allows the sampler to explore negative flux values, but increasingly penalizes them in proportion to how many sigma below zero they fall, thereby favoring physically plausible (positive) fluxes without imposing a hard boundary. Additionally, we include five systematics parameters: a linear slope (), three principal component analysis amplitudes (), and the tilt-event jump amplitude (). We also fit for the Gaussian process amplitude () and timescale (), along with a scatter parameter () that is added in quadrature to the measurement uncertainties.
@@
A Bayesian Information Criterion (BIC) analysis strongly favours the inclusion of the stellar variability model in the Order 2 white light curve (), while there is no preference for its inclusion in the Order 1 white light curve (). A sinusoidal function best describes this variability, leading us to introduce two additional parameters: the stellar variability amplitude () and phase offset (). As a result, the total number of fitted parameters is 21 for the Order 1 white light curve and 23 for Order 2.
@@
For each fit, we initialize the sampler with four walkers per parameter and run for 50,000 steps. We first conduct an initial fit with 12,500 steps, corresponding to 25% of the total run, and use the maximum-probability parameters from this preliminary fit as the starting positions for the final 50,000-step run. To ensure convergence, we discard the first 30,000 steps, or 60% of the total run, as burn-in, retaining only the posteriors from fully converged walkers. We check convergence through visual inspection of trace plots and corner plots, as well as by computing the integrated autocorrelation time (Foreman-Mackey et al., 2013). For both orders, the retained number of steps per walker exceeded 50 times the autocorrelation length.
@@
We fit our spectral light curves by binning the data by pixel with 5 pixels per bin and again normalize the light curves to the first eclipse. For Order 2, we exclude wavelengths longer than m to prevent overlap with Order 1. This results in 113 spectral light curves for Order 2 and 409 for Order 1.
@@
We fix the four orbital parameters to the median values obtained from the posteriors of white light curve analysis: MJD for Order 1, MJD for Order 2, days, and . Systematics parameters are assigned Gaussian priors centered on their best-fit values from the white light curve fit, with standard deviations set to the corresponding measurement uncertainties. Rather than fitting a separate GP model at each wavelength, we scale the mean GP model from the white light curve fit by introducing a scaling factor, (e.g., Radica et al., 2024; Coulombe et al., 2025). Additionally, we fix the projected area of the planet, determined from , based on the fit of Order 1 white light curve (). This value is consistent within 1 of the result from Delrez et al. (2016) (). For the stellar variability model in Order 2, we fix the phase offset to the best-fit value from the white light curve fit and constrain the amplitude using a Gaussian prior informed by the white light curve results.
@@
We adopt the same fitting strategy as for the white light curve. The first spectroscopic bin is fit using 50,000 iterations. For subsequent wavelength bins, the best-fit parameters from the previous bin are used as initial conditions, allowing the sampler to converge more efficiently. These subsequent bins are ran for 12,500 iterations, as they begin closer to their optimal solutions. As in the white light curve fit, each spectroscopic bin is first run for an initial burn-in phase consisting of 25% of the total iterations, after which the fit is restarted at its maximum-likelihood parameters for the full iteration count. To ensure robust posterior sampling, we discard the first 60% of steps as burn-in before analyzing the final parameter distributions.
In both our fits for the white light curve and spectroscopic light curve fits we allow the astrophysical model to fit for negative planetary flux but disfavor such solutions by applying a half-Gaussian prior as described above.
@@
| aBourrier et al. (2020a), bDelrez et al. (2016) | ||||
| †Additional prior on of from Delrez et al. (2016) | ||||
2.3.5 Differences between exoTEDRF and NAMELESS lightcurve fits
Here, we have described the reduction and light curve fitting process used in our analysis with exoTEDRF. As mentioned earlier, an independent reduction and light curve analysis using NAMELESS is detailed in Allart et al. (Under Review) and Pelletier et al. (2025a), with a brief summary provided in Appendix A. However, we provide a brief overview of key differences in the fitting approaches for comparison later in this work.
@@ 2.3.5 Differences between exoTEDRF and NAMELESS lightcurve fits
The exoTEDRF analysis includes ellipsoidal variation in the light curve model, whereas NAMELESS does not.
@@
These differences have several implications. While the contribution of ellipsoidal variation to the total flux is found to be small, neglecting it can lead to a slight overestimation of flux at orbital quadrature. The handling of negative flux presents a more complex challenge. On one hand, negative planetary flux is physically unrealistic, requiring systematic trends to compensate for observed dips. On the other hand, strictly enforcing positive planetary flux risks misrepresenting the true shape of the phase curve. By applying a half-gaussian penalty on negative planetary flux, we aim to bridge the gap between the two approaches. The effect of these fundamental fitting differences are seen most prominently when estimating the nightside flux (see Figures A5). Further, as a negative flux cannot be converted into a brightness temperature, the nightside temperature will have a larger uncertainty (Figure A6).
@@ 2.4 Slice Model
The NIRISS/SOSS wavelength coverage extends into both the optical and near-infrared, allowing us to estimate the planet’s reflected light. If we assume the planet has a gray albedo and emits as a blackbody, a sufficiently high geometric albedo ( 0.2) would result in reflected light dominating at lower wavelengths, likely within Order 2 (see Fig A9). To distinguish reflected light from thermal emission, we decompose the de-trended, phase-dependent planetary flux into these two components:
We model these components using a longitudinal slice approach, following methods applied in previous studies (Cowan & Agol, 2008; Coulombe et al., 2025). The planet is divided into longitudinal slices, where each slice is characterized by a thermal emission component, , and an apparent albedo, . The total planetary flux is then expressed as:
@@
Here, and represent the normalized visibility and illumination of a given planetary location at time (Cowan et al., 2013).
@@
We define the planetary longitude and latitude such that the sub-stellar position is at . The longitudes seen by the observer change with time, , while the observed latitude is equal to the inclination () which we approximate as edge-on, .
@@
To address this, we follow a similar approach to our sinusoidal fits using emcee, but we increase the total number of steps to 100,000 and use 100 walkers. Naïvely, the fit would include parameters: for the albedo values, for the emission parameters, and one additional scatter parameter, . However, since night-side slices do not contribute to the reflected light component, we exclude these albedo values from the fit. In any case, our choice of 100 walkers ensures a sufficient number of walkers per free parameter. Following Coulombe et al. (2025) we set an upper prior limit of on all albedo slices as a fully Lambertian sphere () corresponds to a geometric albedo of . For thermal emission we impose a uniform prior between 0 and 500 ppm for each slice.
We choose to fit our detrended lightcurves considering 4, 6 and 8 longitudinal slices (). However, we show the results of the simplest 4 slice model. As in our previous fits, we conduct an initial run with 25,000 steps (25% of the total run) and use the maximum-probability parameters from this preliminary fit as the starting positions for the final 75,000-step run. We then discard the first 60% of the final run as burn-in.
@@ 2.5 Planetary Effective Temperature
Phase curves are the only way to probe thermal emission from the day and nightside of an exoplanet and hence determine its global energy budget (Parmentier & Crossfield, 2018). The wavelength range of NIRISS/SOSS covers a large portion of the emitted flux of WASP-121 b ( 50–83%; see Figure 2), enabling a precise and robust constraint of the planet’s energy budget.
@@
There are many ways of converting brightness temperatures to effective temperature, including the Error-Weighted Mean (EWM), Power-Weighted mean (PWM) and with a Gaussian Process (Schwartz & Cowan, 2015; Pass et al., 2019). In this work, we elect to compute our effective temperature estimates with a novel method that is essentially a combination of the PWM and EWM. We create the effective temperature by using a simple Monte Carlo process. First, we perturb our emission spectra at each point in the orbit by a Gaussian based on the measurement uncertainty. Our new emission spectrum is then used to create an estimate of the brightness temperature spectrum. This process is repeated at each orbital phase. We then estimate the effective temperature, for a given orbital phase as
@@
We run the Monte Carlo simulations times, propagating all relevant astrophysical uncertainties. Specifically, we propagate the uncertainties of and as determined by the Order 1 white light curve fit (, ). Additionally, we account for the uncertainty in the stellar effective temperature, adopting the value reported in Sing et al. (2024) ( K). We record the median and 1 interval of the effective temperature at all orbital phases.
@@ 3.1 Bond Albedo and Heat Recirculation Efficiency
Using the values for the dayside (K) and nightside (K) hemispheres (where we have inflated the uncertainties by ), we infer solutions of the Bond albedo and the day-night heat recirculation efficiency following Cowan & Agol (2011a):
@@
Following Schwartz & Cowan (2015), we generate a 2000 2000 grid of day- and nightside temperatures as a function of the Bond albedo and heat recirculation efficicency. We then create a surface from equation 20 by comparing these models to our measured dayside and nightside effective temperatures. In order to properly account for the uncertainty in irradiation temperature we recorded and during each iteration of the Monte Carlo simulation from section 2.5. We then find the , , and confidence intervals defined when respectively. For our fit, we infer values of and from temperatures with uncertainties inflated by .
@@ 3.2 Phase Offsets and Relative Amplitude
We use our light curve fits to estimate the phase offset and the phase curve relative amplitude. The phase offset is defined as the orbital phase of the brightest location of the light curve relative to the sub-stellar point. To first order, it is the longitude of the hot-spot (Schwartz et al., 2017). Positive phase offsets correspond to an eastward shift (i.e., maximum flux occurs before mid-eclipse), while negative offsets indicate a westward location. Meanwhile, the phase curve relative amplitude is the normalized amplitude between the maximum flux and the minimum flux (i.e., ). This metric serves as a proxy for the day–night temperature contrast; for example, a relative amplitude of 1 would indicate that one hemisphere emits no detectable flux. Such measurements offer critical insight into the energy balance and heat redistribution in the planetary atmosphere (Cowan et al., 2012; Schwartz & Cowan, 2015).
@@
We compare the relative phase amplitude and phase offsets at different wavelengths in Figure 4. Dobbs-Dixon & Cowan (2017) used a similar figure to show that hot Jupiters observed with Spitzer exhibit larger relative amplitudes and phase offsets at 3.6 m compared to 4.5 m. In an idealized bolometric light curve, we naively might expect phase offset and amplitude to have a one-to-one relationship, where planets with larger phase amplitudes (stronger day-night temperature contrasts) should exhibit smaller phase offsets (Schwartz et al., 2017). However, more recent work shows that hot Jupiters are diverse and that the phase offsets and amplitudes are not necessarily correlated and can be shaped by different components of the circulation (Roth et al., 2024).
@@
We find phase offsets both shifted eastward in the white light curves as 5.1 1.4∘ for Order 1 and 10.5 9.9∘ for Order 2. From our spectroscopic fits we find that the longer wavelengths tend to exhibit small phase offsets and a range of phase curve relative amplitudes (see Fig. 4).
@@
In contrast, at shorter wavelengths we observe an unexpected trend: phase offsets increasing to the east and retaining large relative amplitudes. The high relative amplitude at these wavelengths is due to the nightside emitting little to no flux, making our measured minimum flux effectively zero. This suggests that the day-night recirculation efficiency is still inefficient at these wavelengths, but the brightest location is shifted consistently eastward. At optical wavelengths, we might expect reflected light to impact the offset; however, reflected light measurements tend to favor a westward offset due to condensates evaporating on the hotter eastern hemisphere (Parmentier et al., 2016; Coulombe et al., 2025; Kennedy et al., 2025). Despite the higher irradiation temperature of WASP-121 b, clouds and refractory condensates can still form on the nightside and may be transported to the western limb (Roman et al., 2021; Mikal-Evans et al., 2023; Komacek et al., 2022). Alternatively, the short-wavelength emission might probe different atmospheric depths where the circulation is baroclinic, producing altitude-dependent hotspot shifts (e.g., Changeat et al., 2024). In that case, the observed eastward offset would not represent a permanent equilibrium state.
@@
The surprising result of large eastern offsets at short wavelengths is supported by our independent fits from NAMELESS (see Fig A7). However, the exact values of the offsets and relative amplitudes depend on the fitting framework and on how negative planetary flux is treated. For example, the exoTEDRF fit shows a smaller phase offset than NAMELESS especially within Order 2, likely due to the inclusion of ellipsoidal variation and for Order 2, the stellar variability model. Excluding the stellar variability model brings the offsets from the two fits into closer agreement. Likewise, when the exoTEDRF fitting framework is applied to the NAMELESS reduction, the offsets and relative amplitudes become consistent across the two pipelines.

### v2

@@ 2.3.3 Gaussian Process
We implement a Gaussian process (GP) model to marginalize our inferences over leftover nonwhite noise. We use the open-source Python package celerite (celerite), and consider the SHOTerm Kernel which represents a stochastically driven, damped harmonic oscillator. This kernel models the power spectral density, , using three input parameters: the signal amplitude , characteristic frequency , and quality factor :
@@
The quality factor is set to , a commonly used value for modeling stellar granulation. We use the scaling relations of Gilliland2011 (equations 5 and 6 of that work) and find an expected granulation signal timescale of 6 minutes and an amplitude of 96 ppm, a level we should be sensitive to. Following radica_2024_muted and Coulombe2025-LTT-NIRISS-PC, we adopt the parameterization of Pereira2019-Granulation and reframe the GP parameters in terms of amplitude, , and timescale, . These are then converted back to and before being passed to celerite.
2.3.4 Light-Curve Fitting
@@
We fit our light-curves using the emcee Python package, which implements a Monte Carlo ensemble sampler with affine-invariant Markov chains (emcee). We construct the white light-curves by summing the flux across all 2048 pixels [0.85–2.85 m] for Order 1 and from [0.6–0.85 m] for Order 2. We then normalize both white light-curves to the first eclipse and are then analyzed separately, with the orbital eccentricity fixed to zero (Delrez2016-discovery; Bourrier2020).
@@
For each order, we fit a total of 21–23 parameters. The orbital parameters include the orbital period (), the semi-major axis relative to the stellar radius (), the mid-transit time (), and the orbital inclination (). The astrophysical parameters consist of the planet-to-star radius ratio (), the substellar planet radius to stellar radius ratio (), the quadratic limb-darkening coefficients () (Kipping2013), and the eclipse depth (). To account for phase variations, we fit four sinusoidal parameters (). To avoid unphysical solutions, we impose a prior on the planetary flux (equation 4). Specifically, for any modeled point that yields negative flux, we assign a half-Gaussian prior centered at zero, with the width equal to the measurement uncertainty of that point. This construction allows the sampler to explore negative flux values, but increasingly penalizes them in proportion to how many sigma below zero they fall, thereby favoring physically plausible (positive) fluxes without imposing a hard boundary. Additionally, we include five systematics parameters: a linear slope (), three principal component analysis amplitudes (), and the tilt-event jump amplitude (). We also fit for the Gaussian process amplitude () and timescale (), along with a scatter parameter () that is added in quadrature to the measurement uncertainties.
@@
A Bayesian information criterion (BIC) analysis strongly favors the inclusion of the stellar variability model in the Order 2 white light-curve (), while there is no preference for its inclusion in the Order 1 white light-curve (). A sinusoidal function best describes this variability, leading us to introduce two additional parameters: the stellar variability amplitude () and phase offset (). As a result, the total number of fitted parameters is 21 for the Order 1 white light-curve and 23 for Order 2.
@@
For each fit, we initialize the sampler with four walkers per parameter and run for 50,000 steps. We first conduct an initial fit with 12,500 steps, corresponding to 25% of the total run, and use the maximum-probability parameters from this preliminary fit as the starting positions for the final 50,000-step run. To ensure convergence, we discard the first 30,000 steps, or 60% of the total run, as burn-in, retaining only the posteriors from fully converged walkers. We check convergence through visual inspection of trace plots and corner plots, as well as by computing the integrated autocorrelation time (emcee). For both orders, the retained number of steps per walker exceeded 50 times the autocorrelation length.
@@
We fit our spectral light-curves by binning the data by pixel with 5 pixels per bin and again normalize the light-curves to the first eclipse. For Order 2, we exclude wavelengths longer than m to prevent overlap with Order 1. This results in 113 spectral light-curves for Order 2 and 409 for Order 1.
@@
We fix the four orbital parameters to the median values obtained from the posteriors of white light-curve analysis: MJD for Order 1, MJD for Order 2, days, and . Systematics parameters are assigned Gaussian priors centered on their best-fit values from the white light-curve fit, with standard deviations set to the corresponding measurement uncertainties. Rather than fitting a separate GP model at each wavelength, we scale the mean GP model from the white light-curve fit by introducing a scaling factor, (e.g., radica_2024_muted; Coulombe2025-LTT-NIRISS-PC). Additionally, we fix the projected area of the planet, determined from , based on the fit of Order 1 white light-curve (). This value is consistent within 1 of the result from Delrez2016-discovery (). For the stellar variability model in Order 2, we fix the phase offset to the best-fit value from the white light-curve fit and constrain the amplitude using a Gaussian prior informed by the white light-curve results.
@@
We adopt the same fitting strategy as for the white light-curve. The first spectroscopic bin is fit using 50,000 iterations. For subsequent wavelength bins, the best-fit parameters from the previous bin are used as initial conditions, allowing the sampler to converge more efficiently. These subsequent bins are ran for 12,500 iterations, as they begin closer to their optimal solutions. As in the white light-curve fit, each spectroscopic bin is first run for an initial burn-in phase consisting of 25% of the total iterations, after which the fit is restarted at its maximum-likelihood parameters for the full iteration count. To ensure robust posterior sampling, we discard the first 60% of steps as burn-in before analyzing the final parameter distributions.
In both our fits for the white light-curve and spectroscopic light-curve fits we allow the astrophysical model to fit for negative planetary flux but disfavor such solutions by applying a half-Gaussian prior as described above.
@@
| aBourrier2020, bDelrez2016-discovery | ||||
| †Additional prior on of from Delrez2016-discovery | ||||
2.3.5 Differences between exoTEDRF and NAMELESS light-curve fits
Here, we have described the reduction and light-curve fitting process used in our analysis with exoTEDRF. As mentioned earlier, an independent reduction and light-curve analysis using NAMELESS is detailed in Allart2025 and Pelletier2025, with a brief summary provided in Appendix A. However, we provide a brief overview of key differences in the fitting approaches for comparison later in this work.
@@ 2.3.5 Differences between exoTEDRF and NAMELESS light-curve fits
The exoTEDRF analysis includes ellipsoidal variation in the light-curve model, whereas NAMELESS does not.
@@
These differences have several implications. While the contribution of ellipsoidal variation to the total flux is found to be small, neglecting it can lead to a slight overestimation of flux at orbital quadrature. The handling of negative flux presents a more complex challenge. On the one hand, negative planetary flux is physically unrealistic, requiring systematic trends to compensate for observed dips. On the other hand, strictly enforcing positive planetary flux risks misrepresenting the true shape of the phase curve. By applying a half-gaussian penalty on negative planetary flux, we aim to bridge the gap between the two approaches. The effects of these fundamental fitting differences are seen most prominently when estimating the nightside flux (see Figures C5). Further, as a negative flux cannot be converted into a brightness temperature, the nightside temperature will have a larger uncertainty (Figure C6).
@@ 2.4 Slice Model
The NIRISS/SOSS wavelength coverage extends into both the optical and near-infrared, allowing us to estimate the planet’s reflected light. If we assume the planet has a gray albedo and emits as a blackbody, a sufficiently high geometric albedo ( 0.2) would result in reflected light dominating at lower wavelengths, likely within Order 2 (see Fig C9). To distinguish reflected light from thermal emission, we decompose the detrended, phase-dependent planetary flux into these two components:
We model these components using a longitudinal slice approach, following methods applied in previous studies (Cowan&Agol2008; Coulombe2025-LTT-NIRISS-PC). The planet is divided into longitudinal slices, where each slice is characterized by a thermal emission component, , and an apparent albedo, . The total planetary flux is then expressed as:
@@
Here, and represent the normalized visibility and illumination of a given planetary location at time (Cowan2013).
@@
We define the planetary longitude and latitude such that the substellar position is at . The longitudes seen by the observer change with time, , while the observed latitude is equal to the inclination () which we approximate as edge-on, .
@@
To address this, we follow a similar approach to our sinusoidal fits using emcee, but we increase the total number of steps to 100,000 and use 100 walkers. Naïvely, the fit would include parameters: for the albedo values, for the emission parameters, and one additional scatter parameter, . However, since nightside slices do not contribute to the reflected light component, we exclude these albedo values from the fit. In any case, our choice of 100 walkers ensures a sufficient number of walkers per free parameter. Following Coulombe2025-LTT-NIRISS-PC we set an upper prior limit of on all albedo slices as a fully Lambertian sphere () corresponds to a geometric albedo of . For thermal emission we impose a uniform prior between 0 and 500 ppm for each slice.
We choose to fit our detrended light curves considering four, six and eight longitudinal slices (). However, we show the results of the simplest four slice model. As in our previous fits, we conduct an initial run with 25,000 steps (25% of the total run) and use the maximum-probability parameters from this preliminary fit as the starting positions for the final 75,000-step run. We then discard the first 60% of the final run as burn-in.
@@ 2.5 Planetary Effective Temperature
Phase curves are the only way to probe thermal emission from the day- and nightside of an exoplanet and hence determine its global energy budget (Phasecurve-book). The wavelength range of NIRISS/SOSS covers a large portion of the emitted flux of WASP-121 b ( 50–83%; see Figure 2), enabling a precise and robust constraint of the planet’s energy budget.
@@
There are many ways of converting brightness temperatures to effective temperature, including the Error-Weighted Mean (EWM), Power-Weighted mean (PWM) and with a Gaussian process (Schwartz2015; Pass2019-TeffGP). In this work, we elect to compute our effective temperature estimates with a novel method that is essentially a combination of the PWM and EWM. We create the effective temperature by using a simple Monte Carlo process. First, we perturb our emission spectra at each point in the orbit by a Gaussian based on the measurement uncertainty. Our new emission spectrum is then used to create an estimate of the brightness temperature spectrum. This process is repeated at each orbital phase. We then estimate the effective temperature, for a given orbital phase as
@@
We run the Monte Carlo simulations times, propagating all relevant astrophysical uncertainties. Specifically, we propagate the uncertainties of and as determined by the Order 1 white light-curve fit (, ). Additionally, we account for the uncertainty in the stellar effective temperature, adopting the value reported in Sing2024-NIRSpec ( K). We record the median and 1 interval of the effective temperature at all orbital phases.
@@ 3.1 Bond Albedo and Heat Recirculation Efficiency
Using the values for the dayside (K) and nightside (K) hemispheres (where we have inflated the uncertainties by ), we infer solutions of the Bond albedo and the day-night heat recirculation efficiency following Cowan2011-albedo:
@@
Following Schwartz2015, we generate a 2000 2000 grid of day- and nightside temperatures as a function of the Bond albedo and heat recirculation efficicency. We then create a surface from equation 20 by comparing these models to our measured dayside and nightside effective temperatures. In order to properly account for the uncertainty in irradiation temperature we recorded and during each iteration of the Monte Carlo simulation from section 2.5. We then find the , , and confidence intervals defined when respectively. For our fit, we infer values of and from temperatures with uncertainties inflated by .
@@ 3.2 Phase Offsets and Relative Amplitude
We use our light-curve fits to estimate the phase offset and the phase curve relative amplitude. The phase offset is defined as the orbital phase of the brightest location of the light-curve relative to the substellar point. To first order, it is the longitude of the hot spot (Schwartz2017). Positive phase offsets correspond to an eastward shift (i.e., maximum flux occurs before mid-eclipse), while negative offsets indicate a westward location. Meanwhile, the phase curve relative amplitude is the normalized amplitude between the maximum flux and the minimum flux (i.e., ). This metric serves as a proxy for the day–night temperature contrast; for example, a relative amplitude of 1 would indicate that one hemisphere emits no detectable flux. Such measurements offer critical insight into the energy balance and heat redistribution in the planetary atmosphere (Cowan2012; Schwartz2015).
@@
We compare the relative phase amplitude and phase offsets at different wavelengths in Figure 4. Dobbs-Dixon&Cowan2017 used a similar figure to show that hot Jupiters observed with Spitzer exhibit larger relative amplitudes and phase offsets at 3.6 m compared to 4.5 m. In an idealized bolometric light-curve, we naively might expect phase offset and amplitude to have a one-to-one relationship, where planets with larger phase amplitudes (stronger day-night temperature contrasts) should exhibit smaller phase offsets (Schwartz2017). However, more recent work shows that hot Jupiters are diverse and that the phase offsets and amplitudes are not necessarily correlated and can be shaped by different components of the circulation (Roth2024).
@@
We find phase offsets both shifted eastward in the white light-curves as 5.1 1.4∘ for Order 1 and 10.5 9.9∘ for Order 2. From our spectroscopic fits we find that the longer wavelengths tend to exhibit small phase offsets and a range of phase curve relative amplitudes (see Fig. 4).
@@
In contrast, at shorter wavelengths we observe an unexpected trend: phase offsets increasing to the east and retaining large relative amplitudes. The high relative amplitude at these wavelengths is due to the nightside emitting little to no flux, making our measured minimum flux effectively zero. This suggests that the day-night recirculation efficiency is still inefficient at these wavelengths, but the brightest location is shifted consistently eastward. At optical wavelengths, we might expect reflected light to impact the offset; however, reflected light measurements tend to favor a westward offset due to condensates evaporating on the hotter eastern hemisphere (Parmentier2016; Coulombe2025-LTT-NIRISS-PC; Kennedy2025). Despite the higher irradiation temperature of WASP-121 b, clouds and refractory condensates can still form on the nightside and may be transported to the western limb (Roman2021; Mikal-Evans2023; Komacek-ultrahotclouds). Alternatively, the short-wavelength emission might probe different atmospheric depths where the circulation is baroclinic, producing altitude-dependent hot spot shifts (e.g., Changeat2024). In that case, the observed eastward offset would not represent a permanent equilibrium state.
@@
The surprising result of large eastern offsets at short wavelengths is supported by our independent fits from NAMELESS (see Fig C7). However, the exact values of the offsets and relative amplitudes depend on the fitting framework and on how negative planetary flux is treated. For example, the exoTEDRF fit shows a smaller phase offset than NAMELESS especially within Order 2, likely due to the inclusion of ellipsoidal variation and for Order 2, the stellar variability model. Excluding the stellar variability model brings the offsets from the two fits into closer agreement. Likewise,
when the exoTEDRF fitting framework is applied to the NAMELESS reduction, the offsets and relative amplitudes become consistent across the two pipelines.